The waiting time for a train has a uniform distribution between 3 and 22 minutes. What is the probability that the waiting time for this train is more than 5 minutes on a given day? Answer: (Round to four decimal places.)
Answer #1

Solution :

Given that ,

a = 3

b = 22

The probability that the waiting time for this train is more than 5 minutes on a given day,

P( X > c ) = ( b - c ) / ( b - a )

P( X > 5 ) = ( 22 - 5 ) / ( 22 - 3 )

= 17 / 19

= 0.8947